# AI-Enhanced Self-Triggering for Extensive Air Showers: Performance and FPGA Feasibility

### v1

@@ 4.1.4 Example of a Low-Amplitude Signal Classification
5 Estimated Latency on FPGA Platforms
To evaluate the feasibility of real-time deployment, the model’s inference latency was first measured on an Apple Silicon GPU using Metal Performance Shaders (MPS). Each inference took approximately per trace, which implies a required buffer of around 8,250 samples when sampling at . Since real-time streaming at this rate is continuous, reducing inference latency is essential to minimize buffering and avoid potential pipeline stalls.
Moving from a GPU to an FPGA requires a different set of assumptions compared to traditional CPU-to-FPGA speed-up estimates. In particular, FPGAs allow for specialized parallelism and hardware-friendly dataflow optimizations that can significantly reduce latency if designed properly. Table 1 summarizes conservative latency estimates for different FPGA classes, along with the corresponding buffer sizes at . These values were derived by adjusting for GPU-to-FPGA acceleration, assuming an optimized hardware design that capitalizes on the inherent parallelism of FPGAs.
FPGA Type / Design Flow | Estimated Latency (per trace) | Required Buffer Size (samples) |
Low-end FPGA (Zynq) | 11.17 s | 2,792 samples |
Mid-range FPGA (Arria 10) | 4.78 s | 1,195 samples |
High-end FPGA (Alveo U280) | 2.23 s | 558 samples |
High-end (INT8-Optimized) | 1.12 s | 280 samples |
Here, the row labeled High-end (INT8-Optimized) represents a design flow that leverages 8-bit quantization on a device similar to other high-end FPGAs, rather than indicating a distinct family of boards. Narrower data paths in an INT8 pipeline can further reduce latency and resource utilization, at the expense of additional quantization efforts.
Even a low-end FPGA such as a Zynq device could reduce the baseline GPU latency from down to approximately , significantly lowering the buffering requirement. However, whether this meets “real-time constraints” depends on the total system requirements, including any overhead from I/O, external memory transfers, and other pipeline stages. In large-scale or power-sensitive deployments, factors such as cost, power consumption, and design complexity also guide the choice of FPGA platform and quantization scheme. The final selection must balance these considerations to ensure both feasibility and efficiency for the target application.
To fully enable real-time classification of cosmic-ray radio pulses, a dedicated FPGA-based framework must balance latency, resource utilization, and quantization precision while integrating seamlessly with experimental data streams. The results in Table 1 demonstrate that even low-end FPGA platforms can achieve substantial latency reductions, but practical deployment also requires optimizing DSP and memory utilization to handle high-rate data streams efficiently. Implementing 8-bit quantization (INT8) can further reduce both inference latency and resource consumption, although additional effort is needed to mitigate potential accuracy loss. A real-time classification system must also account for I/O bandwidth constraints, buffering mechanisms, and event prioritization to avoid pipeline stalls. Future work will focus on refining this framework by incorporating real-time data acquisition, firmware-level optimizations, and adaptive triggering strategies, ensuring a robust and scalable AI-driven self-triggering system for next-generation cosmic-ray experiments.
Future work will extend beyond optimizing the model for FPGA deployment while preserving classification performance. A key focus will be on benchmarking against alternative AI architectures, refining network optimization techniques, and expanding the training dataset to enhance generalization. Incorporating extensive air shower (EAS) simulations with state-of-the-art tools such as CoREAS will provide a more comprehensive and physically motivated dataset. Additionally, integrating experimental data from particle detector-triggered events will create an unprecedentedly realistic training set. Our preliminary results indicate that training on high-amplitude datasets enables the detection of even the faintest radio pulses—signals that have remained undetected by traditional methods to date. By leveraging these advancements, we aim to push the boundaries of real-time cosmic ray detection, making AI-driven self-triggering a transformative technology for next-generation astroparticle physics experiments.
6 Conclusion
In this work, we have demonstrated that artificial intelligence can dramatically enhance the self-triggering capability of radio detectors for extensive air showers (EAS). By training on large volumes of measured noise data and injected cosmic-ray-like pulses, our deep learning pipeline—incorporating both an autoencoder-based denoiser and a fully convolutional classifier—achieves excellent separation of genuine signals from background. Over a wide range of signal-to-noise ratios, the approach retains efficiencies in excess of 99% while simultaneously suppressing false positives down to sub-percent levels. Even in an urban-like noise environment, with RFI sources and transient interference, the model maintains robust performance, indicating strong potential for deployment under realistic conditions. Crucially, the flexibility of our inference stage enables an operating point where false positives can be entirely eliminated (as observed in our large-scale noise dataset), albeit at a modest reduction in true-positive rate to roughly 72%. This result opens the door to single-station cosmic-ray detection purely by radio triggers, removing the need to rely on any particle detectors for trigger generation. Although multiple-station coincidences remain a valuable method for additional false-positive suppression, these findings show that a single antenna unit could be made sufficiently selective to capture genuine air-shower events in high-noise environments. A key outcome of this study is the successful demonstration of latency estimates that make real-time triggering feasible. Our analysis of FPGA implementations shows that when using optimized designs and dataflow pipelines, inference times can reach the microsecond regime. This aligns well with the strict timing requirements of large-scale radio observatories, where trigger decisions must be made rapidly to record short-lived air shower signals. The ability to process data in-stream, directly on reconfigurable hardware, underscores the practical viability of these AI-driven methods for next-generation detectors. Overall, these results address one of the longstanding challenges in radio-based cosmic ray detection: reducing or eliminating the need for external triggers from particle or fluorescence detectors. An autonomous trigger operating purely on radio data lowers complexity and costs, while also unlocking sensitivity to high zenith angles and potentially exotic UHE particle interactions. Moving forward, further integration with simulation tools (e.g. CoREAS) and additional real shower data will refine the models’ accuracy and increase robustness to a wider spectrum of signal morphologies. The high efficiency and low false trigger rates shown here pave the way for a new class of AI-based, fully self-triggered radio arrays, promising both improved coverage and enhanced scientific reach in ultra-high-energy cosmic ray research.
7 Discussion
The results presented in this study highlight several key insights into the performance and robustness of our AI-driven self-triggering framework for radio detectors. Beyond achieving high detection efficiencies and minimal false-positive rates, a deeper examination of the model’s behaviour reveals its ability to generalise well across different noise conditions, preserve signal fidelity through denoising, and maintain robustness in realistic deployment scenarios.
A notable finding is the classifier’s ability to maintain high purity on non-denoised background traces, even though it was never trained on them. Because the classifier was trained exclusively on denoised signal and background data, one might have expected it to misinterpret raw background as containing signals. However, testing showed that it continued to distinguish signal from background with high accuracy, suggesting it was not misled by potential artifacts introduced by the denoiser on background samples. This indicates that the combined denoiser–classifier pipeline does not merely adapt to denoiser-induced modifications but instead learns robust, physically meaningful pulse features, demonstrating strong generalization across different noise conditions. Furthermore, these results suggest that mixing raw and denoised data in the training set may be an intriguing avenue for future exploration, as it could further improve robustness to a wider range of conditions.
We also investigated the effect of narrow-band RFI suppression. For the datasets discussed in this work, we did not initially remove strong RFI components. However, we produced an alternate noise dataset in which these narrow-band RFI components were suppressed prior to pulse injection. Because suppressing RFI lowers the noise floor by about 17 dB, we scaled down the injected pulses by a matching factor to preserve the same relative signal-to-noise ratio (SNR) distribution that our model was originally trained on. As expected, the classifier’s performance remained virtually unchanged, since the effective classification task difficulty was not altered. Nevertheless, eliminating RFI could simplify the denoiser and classifier in a real-world deployment, as it reduces background complexity. A simpler network architecture may lead to faster inference times and lower computational costs, paving the way for real-time triggering with reduced latency.
Another key consideration for real-world applications is the resilience of our pipeline to variations in noise characteristics. Our tests in an urban-like interference environment suggest that the denoiser–classifier chain is learning fundamental signal-versus-noise traits rather than overfitting to specific noise profiles. Further studies could systematically probe this robustness using augmented training datasets (e.g., artificially injecting varied noise profiles) or by testing on independent, geographically diverse datasets.
Beyond classification accuracy, latency remains critical for real-time self-triggering applications. Preliminary estimates indicate that microsecond inference is feasible on FPGA platforms, but hardware-specific optimisations are needed to confirm these figures in practice. Refining the neural network architectures to strike an optimal balance between complexity and speed could ensure that real-time decision-making constraints are met without compromising accuracy.
Looking ahead, our framework could be extended by incorporating CoREAS-simulated pulses covering a wide range of air shower scenarios, including variations in energy, zenith angle, and primary composition. Training on these simulated signals would allow the model to learn a more comprehensive representation of shower-induced radio emission, likely improving its ability to detect real cosmic-ray pulses in challenging regimes. This approach would also provide a clear path for systematically evaluating the pipeline’s performance under diverse air shower conditions.
Finally, although our results demonstrate the feasibility of single-antenna triggering, integrating AI-driven self-triggering with multi-station coincidence algorithms is an exciting possibility. Cross-station coincidence can lower false-positive rates while preserving sensitivity to rare events, including highly inclined cosmic rays that might elude traditional triggers.
In summary, our findings underscore the transformative potential of deep learning for self-triggered radio detection of extensive air showers. By preserving robust generalisation, demonstrating strong performance even with various noise conditions, and operating at latencies amenable to real-time applications, AI-based triggering stands poised to play an integral role in next-generation cosmic-ray observatories. Future work will focus on refining model architectures, evaluating performance across different noise environments, integrating large-scale radio arrays, and exploring how single-station AI triggers may combine with multi-station coincidence for enhanced detection confidence.
@@ Acknowledgments
This research made use of open-source software, including PyTorch and various scientific computing libraries in Python, such as NumPy, SciPy, and Matplotlib. The author acknowledges their developers for providing these valuable tools, which played a crucial role in data processing, model training, and analysis.

### v3

@@ 4.2 Example of a Low-Amplitude Signal Classification
4.3 Comparison with a Threshold-Based Trigger
To provide a baseline against conventional self-triggering techniques, I implemented a simple SNR-threshold trigger. For each evaluation trace, I form the analytic signal via the Hilbert transform and define the score as in Eq. 2.1. For background traces, the trace itself serves as this reference, so the largest envelope excursion is treated as a spurious candidate signal. The resulting distributions of scores for signal and background events allow the construction of a ROC curve by varying the threshold on the SNR score.
To ensure a fair comparison, both the neural-network classifier and the SNR-threshold baseline were evaluated on the same unfiltered datasets, which contain the full interference environment as described in Section 2.1. In practice, threshold-based self-triggers are often preceded by frequency filtering to increase sensitivity, but such preprocessing would also be expected to benefit the neural network if applied consistently. Applying additional filtering only to the baseline would artificially simplify its input and alter the effective noise conditions relative to the classifier evaluation. For this reason, I restrict the comparison to identical unfiltered inputs for both methods.
For a fixed input, ROC performance is invariant under any monotonic reparameterisation of the score. For example, using the peak-to-RMS ratio or the z-score is mathematically equivalent to our SNR definition up to a monotonic transform and therefore yields the same ROC curve within numerical precision. By contrast, using the raw peak voltage without noise normalisation performs significantly worse, as expected, because it does not account for fluctuations in the noise level. More elaborate variants (e.g. running-average baselines, short-time energy) remain threshold triggers in spirit but introduce method-specific hyper-parameters and design choices. To avoid cherry-picking, I therefore restrict the comparison to this minimal representative baseline.
Figure 5 shows the ROC of the SNR-threshold trigger. The area under the curve is . At the operating point used throughout this work, corresponding to a false positive rate of (vertical red dashed line), the threshold-based trigger achieves a signal efficiency of only . For comparison, the neural-network classifier maintains an efficiency of under the same conditions (cf. Fig. 3). This demonstrates a clear improvement of the AI-based approach over a traditional threshold trigger in the challenging high-interference environment considered here.
4.4 Physics Implications of Improved Triggering
Beyond the technical demonstration, it is instructive to consider the physics impact of lowering the trigger threshold in realistic, interference-rich environments. At the chosen operating point (false positive rate of ), the SNR-threshold baseline achieves a signal efficiency of only (Sec. 4.3), whereas the neural-network classifier retains an efficiency of (Sec. 4). This corresponds to an effective reduction of the detectable signal-to-noise threshold by more than a factor of two.
Such an improvement has direct consequences for the reach of radio arrays. Lowering the trigger threshold extends the accessible energy range towards lower primary energies, where air-shower radio signals are intrinsically weaker. It also improves sensitivity to very inclined air showers, since the particle component is largely absorbed in the atmosphere while the radio emission remains strong. These events are particularly valuable for composition studies, for enlarging the sky coverage of ground-based observatories, and as a prerequisite for detecting ultra-high-energy neutrinos, which manifest as upward-going or near-horizontal showers. By enabling robust radio-only operation under high-background conditions, the method reduces reliance on external particle detectors, thereby simplifying array design and allowing sparser station layouts. This opens a path toward larger effective apertures at EeV energies, complementing and extending the capabilities of existing cosmic-ray and neutrino observatories.
These considerations show that the proposed AI-based trigger is not only technically feasible on FPGA hardware, but also provides a tangible physics benefit by increasing the accessible parameter space of radio arrays in energy, inclination, and coverage.
5 Quantisation and FPGA Implementation
5.1 Quantisation with HLS4ML
For real-time deployment, the trained floating-point classifier was converted to a fixed-point representation using the HLS4ML framework [23, 24, 25]. Quantisation replaces floating-point multiplications and additions with fixed-point arithmetic, reducing resource usage and latency on FPGAs while maintaining compatibility with hardware synthesis workflows. The key challenge is to choose a precision that minimises resource cost without sacrificing classification performance.
To determine a suitable fixed-point format for FPGA deployment, I first estimated the minimum required integer bit-width to avoid overflow. This was done by forwarding representative validation samples through the trained model while recording the maximum absolute activation values across all intermediate layers. Based on this analysis, I fixed the integer width at 5 bits.
Subsequently, I scanned the total bit-width while keeping the integer part fixed, evaluating classification performance at each precision setting. Figure 6a shows the area under the ROC curve (AUC) as a function of total bit-width. Performance remains stable down to 13 total bits, below which degradation becomes apparent. I therefore selected ap_fixed<13,5> (13 total bits, 5 integer bits) as the default precision for FPGA synthesis.
5.2 Float vs. Quantised Model Performance
The effect of quantisation on classification performance was evaluated by comparing the floating-point and ap_fixed<13,5> implementations on the same validation datasets.
As shown in Figure 6b, the ROC curves for the two models are nearly identical, with the AUC decreasing only marginally from the PyTorch floating-point baseline () to the quantised implementation ().
This confirms that the chosen fixed-point precision preserves the model’s separation power between signal and background, enabling an efficient FPGA implementation without loss of physics performance.
5.3 FPGA Synthesis with Vitis 2023
The quantised model was synthesised using Vitis HLS 2023.2 [28], targeting three representative FPGA families: the low-cost Zynq-7000 (xc7z020), the mid-range Kintex UltraScale (xcku040), and the high-end Alveo U250 (xcu250). A target clock period of () was assumed in all cases. The measured end-to-end latency, pipeline initiation interval (II), and DSP usage are summarised in Table 1.
| Platform | Latency [s] | II | DSP usage | Timing met |
| Zynq-7000 xc7z020 | 6.51 | 1060 | 963 % | yes |
| Kintex UltraScale xcku040 | 3.30 | 654 | 108 % | yes |
| Alveo U250 xcu250 | 1.99 | 394 | 17 % | yes |
5.3.1 Resource Utilisation
DSP slices are the critical limiting resource for this design, as they are heavily used for multiply–accumulate (MAC) operations in convolutional and dense layers. While LUT and BRAM usage remain well below device limits in all cases, DSP saturation is encountered for the Zynq-7000 and, to a lesser extent, the Kintex UltraScale. This constraint dictates optimisation strategies, such as resource sharing or further model compression, for deployment on cost-sensitive platforms. The Alveo U250, by contrast, offers substantial headroom for scaling up the model or reducing latency further.
5.3.2 Latency
Figure 7 shows the measured inference latency for the three synthesis targets. Even without aggressive optimisation, the high-end Alveo U250 achieves sub- latency. The Kintex UltraScale operates at , while the Zynq-7000 delivers . All results meet the -level latency guideline for station-level triggering in the target application.
5.4 Power consumption
I characterised fabric-only on-chip power for the HLS4ML-generated accelerator implemented on a Xilinx Alveo U250 (xcu250-figd2104-2L-e) using Vivado 2023.2. To measure only the neural-network accelerator’s contribution, I built a minimal ‘fabric-only’ design: the HLS4ML core was the sole processing block at the top level, with no platform shell, memory controllers, PCIe or other peripherals, and no external I/O interfaces. The only additional logic was an on-chip pseudo-random bit sequence (PRBS) generator to provide a continuous input stimulus. The design was driven by a single clock and taken through full implementation (synthesis, placement, and routing) so power was reported on the placed-and-routed netlist.
The PRBS source was implemented as a 32-bit LFSR with taps , replicated across the input bus, with handshake/control pins held active to keep the pipeline filled. This stimulus exercises the full datapath and prevents constant-propagation through the network.
Post-implementation power was obtained with Vivado’s vectorless analysis (no activity annotation). The placed-and-routed netlist reports a dynamic power of for the core. If clock-network power is also excluded, the dynamic component attributable solely to datapath logic is approximately . Because the estimate remains vectorless and excludes platform infrastructure, the reported numbers should be interpreted as a lower bound; absolute power is expected to increase with activity-annotated runs and board-level I/O. As the U250 is the largest of the three target platforms, this value is conservative with respect to device scaling.
These results indicate ample headroom within a typical budget for a compact radio-detection pipeline, leaving margin for front-end filtering, data conversion, and control logic. A full platform-inclusive power profile (including activity-annotated power and comparisons of LUT- vs. DSP-mapped arithmetic) is left for future hardware studies.
6 Summary and Outlook
I have presented an artificial-intelligence-based self-trigger for radio detection of extensive air showers and demonstrated its feasibility for real-time deployment on FPGA hardware. Starting from a floating-point model trained in PyTorch on measured noise from a high-interference environment combined with cosmic-ray pulses generated using the MGMR3D simulation framework, the workflow covers a broad range of shower geometries, energies, and antenna–core distances. The trained fully convolutional classifier achieves stable performance under operating conditions representative of MHz-scale trigger-trial rates encountered in urban or RFI-rich sites.
For context, I also compared the neural-network classifier to a simple threshold-based trigger, implemented as an SNR cut on the Hilbert-envelope peak relative to the noise variance. The baseline achieves only modest separation power () and a signal efficiency of at the target operating point of , whereas the classifier retains efficiency under the same conditions. This comparison demonstrates that the AI-based approach provides a substantial improvement over traditional threshold triggers in the challenging high-interference environment considered here.
The network was quantised using HLS4ML to a fixed-point precision of ap_fixed<13,5>.
A precision scan confirmed that 13-bit fixed-point arithmetic preserves the physics performance of the floating-point model (, ).
The quantised model was synthesised with Vitis HLS 2023.2 for three FPGA platforms (Zynq-7000, Kintex UltraScale, and Alveo U250).
High-level synthesis results show inference latencies in the microsecond range, meeting the first-level trigger timing budget for large-scale radio arrays.
DSP usage emerged as the primary resource bottleneck, guiding optimisation strategies for deployment on cost-constrained devices. A post-implementation, fabric-only power estimate on the Alveo U250 reports at (vectorless analysis), which I interpret as a conservative lower bound on on-chip power.
For the datasets discussed in this work, strong narrow-band RFI components were not removed prior to training. In a real-world deployment, eliminating such RFI using an FFT-based filter implemented on the FPGA could simplify the classification task by reducing background complexity. A simpler network architecture may lead to faster inference times and lower computational costs, paving the way for real-time triggering with further reduced latency and resource usage.
This work demonstrates, for the first time in the context of radio-detector self-triggering, an end-to-end workflow from AI model training to FPGA-oriented deployment using HLS4ML: quantisation to fixed-point, high-level synthesis with Vitis HLS 2023.2, and validation of physics-performance preservation together with post-synthesis latency/resource characterisation and a fabric-only, post-implementation power estimate. The approach reduces reliance on external triggers from particle or fluorescence detectors and supports autonomous operation scenarios for next-generation radio arrays.
Future developments include broadening the training to a wider family of simulated pulse morphologies and incorporating additional measured air-shower events. On the hardware side, it will be important to complement the fabric-only power estimate reported here with activity-annotated and platform-inclusive profiles across the target devices, and to investigate architectural and clocking optimisations suitable for low-power autonomous stations. Together with simple on-FPGA pre-filtering of narrow-band RFI, these directions are expected to further reduce resource usage and latency and to de-risk deployment in operational radio arrays.
In addition to these technical aspects, the comparison with a simple SNR-threshold baseline highlights the physics impact of the proposed method. At the chosen operating point, the neural-network trigger lowers the effective trigger threshold significantly compared to the SNR baseline. This improved threshold increases sensitivity to weak signals and to very inclined air showers, where the particle cascade is absorbed while the radio emission remains detectable. Such events are essential not only for composition studies and for extending the sky coverage of cosmic-ray observatories, but also as a prerequisite for detecting ultra-high-energy neutrinos with ground-based radio arrays. Moreover, the approach is designed to operate robustly in environments with strong and variable radio backgrounds, where conventional threshold triggers fail, thereby enabling autonomous operation under realistic field conditions. These results demonstrate that FPGA-based AI self-triggering not only meets the technical feasibility requirements but also provides a tangible enhancement of the scientific potential of next-generation radio and neutrino observatories.
@@ Acknowledgments
The author also gratefully acknowledges Prof. Katie Mulrey and Prof. Harm Schoorlemmer (Radboud University Nijmegen) for kindly providing the antenna employed in this study. The assistance of Noah Siegemund in setting up the hardware on campus is also gratefully acknowledged.
This research made use of open-source software, including PyTorch, HLS4ML and various scientific computing libraries in Python, such as NumPy, SciPy, and Matplotlib. The author acknowledges their developers for providing these valuable tools, which played a crucial role in data processing, model training, and analysis.